OpenStudy (anonymous):

A couple have bought a new house and are comparing quotes from two moving companies for moving their furniture. Company A charges $90 for the truck and $60 per hour for the movers. Company B charges $70 for the truck and $70 per hour for the movers. Create a cost equation for each company where y is the total cost and x is the number of hours of labor. Write a system of equations.

OpenStudy (anonymous):

A:y=90+60x...............B:y=70+70x............

OpenStudy (anonymous):

THANK you! ... sometimes I feel like there is more to a question & then it ends up being simple :)